ganeshie8 (ganeshie8):

Are there any positive integers that cannot be expressed as sum of at least two consecutive integers ?

ganeshie8 (ganeshie8):

2 = not possible 3 = 1+2 4 = not possible 5 = 2+3 6 = 1+2+3 7 = 3+4 ...

OpenStudy (omarbirjas):

8

ganeshie8 (ganeshie8):

yes 8 also cannot be expressed as sum of consecutive integers

Parth (parthkohli):

Let the first number be \(p\) and let there be \(q\) consecutive integers.\[k = pq + \dfrac{q(q-1)}{2}\]

OpenStudy (omarbirjas):

to tired to do any more math...

Parth (parthkohli):

I don't know how recasting the problem into this form helps, but let's see.

ganeshie8 (ganeshie8):

that looks good because it gives us something to play with

ganeshie8 (ganeshie8):

k = p + (p+1) + (p+2) + ... + (p+q-1) = pq + q(q-1)/2 so k must be of this form is it

Parth (parthkohli):

exactly.

Parth (parthkohli):

Umm, I haven't even answered the question...

ganeshie8 (ganeshie8):

we're 90% done actually

ganeshie8 (ganeshie8):

\[k = pq + \dfrac{q(q-1)}{2}\] multiplying 2 through out gives \[2k = q(2p+q-1)\] clearly the factors on right hand side cannot have same parity (cannot be both odd or both even)

ganeshie8 (ganeshie8):

thats because the sum of factors : \(q+2p+q-1 = 2(p+q)-1\) is odd

ganeshie8 (ganeshie8):

so \(2k=q(2p+q-1)\) is impossible when \(k\) is a power of \(2\)

OpenStudy (anonymous):

can someone help me with trigonometric raios? sin(3x+13) = cos(4x)